OpenStudy (anonymous):

The system of equations is coincident. What are the missing values? x+3y=5 4x+(blank)y=(blank)

OpenStudy (anonymous):

@MissNoName

OpenStudy (carlyleukhardt):

4x+12y=20

OpenStudy (anonymous):

thanks :D can you help with one more?

OpenStudy (carlyleukhardt):

yeah sure :)

OpenStudy (anonymous):

K, JUST A SEC :)

OpenStudy (anonymous):

oops...caps XD

OpenStudy (carlyleukhardt):

XD lol

OpenStudy (anonymous):

i can help too!

OpenStudy (anonymous):

Match each item in Column A to an answer in Column B. What are the classifications of each system? Column A 1. 4x+y=8 x+3y=8 2. -4x+3y=-2 2x-3y=1 3. 5x-2y=4 5x-2y=6 4. y=3x-4 -3x+y=2

OpenStudy (anonymous):

A. consistent independent B. coincident C. inconsistent

OpenStudy (anonymous):

hey Miss, can you give Carly a medal so I can give you one? XD

OpenStudy (anonymous):

I have to match what those are with A. consistent independent B. coincident C. inconsistent

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

Match each item in Column A to an answer in Column B. What are the classifications of each system? Column A 1. 4x+y=8 x+3y=8 2. -4x+3y=-2 2x-3y=1 3. 5x-2y=4 5x-2y=6 4. y=3x-4 -3x+y=2 ___________________________ column B A. consistent independent B. coincident C. inconsistent

OpenStudy (anonymous):

do you know this one?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@paki XD

OpenStudy (anonymous):

A goes with what? number 1, 2, 3, or 4? XD

OpenStudy (paki):

you know consistent and inconsistent...?

OpenStudy (anonymous):

not really...

OpenStudy (anonymous):

@adelyn_helps

OpenStudy (anonymous):

@MissNoName do you have any idea's on this...? XD

OpenStudy (adelyn_helps):

ok i here :D

OpenStudy (anonymous):

:D

OpenStudy (adelyn_helps):

hang on figuring this out

OpenStudy (anonymous):

its probably A

OpenStudy (anonymous):

XD

OpenStudy (paki):

lol

OpenStudy (anonymous):

I found dah answers! yay :D

OpenStudy (anonymous):

O_O and I got a 100% ...best test time ever...

OpenStudy (adelyn_helps):

ok sorry Im a little slow

OpenStudy (anonymous):

For the graph shown, select the statement that best represents the given system of equations. https://static.k12.com/calms_media/media/316000_316500/316443/1/b5ae79411710c81d3a514f2276d2d71e7e647413/MS_ALG_S1_08_01_quiz_Q2_question.gif 2y – x = 2 x – 2y = 8 -->A. inconsistent B. consistent and independent C. coincident D. not enough information the answer is A.

OpenStudy (anonymous):

uh

OpenStudy (anonymous):

For the graph shown, select the statement that best represents the given system of equations. https://static.k12.com/calms_media/media/1312500_1313000/1312950/1/7a688636620d17ab1b7fedff1dd16087325d6c12/HS_A1G2_S1_08_01_ART3.jpg 3y – x = 12 x – 3y = 6 A. coincident B. consistent and independent --> C. inconsistent D. not enough information this one is C

OpenStudy (anonymous):

How many solutions does this system of equations have? https://static.k12.com/calms_media/media/1513500_1514000/1513629/1/fd224fb15d72187e706697db35a4853f8835a8d4/MS_IMC-1400804-100521.jpg -->A. none B. exactly one C. exactly two D. infinitely many this one is A

OpenStudy (anonymous):

WHAT, okay why is

OpenStudy (anonymous):

im posting the aswers @MissNoName this is for others XD

OpenStudy (adelyn_helps):

I dont no this one sorry I have to go have more school

OpenStudy (anonymous):

WHAT, eh..okay

OpenStudy (anonymous):

The system of equations is coincident. What are the missing values? x+3y=5 4x+(12)y=(20) missing values=12 and 20

OpenStudy (anonymous):

the last one was A B C C

OpenStudy (adelyn_helps):

good luck